The Crash Davis Driving School has an ROE of 16.3 percent and a payout ratio of 62 percent. What is its sustainable growth rate? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Sustainable growth rate %
Answer #1

Answer - 6.60%

Reason -

Sustainable growth rate = [(ROE)(b)] / [1 – (ROE)(b)]

where, ROE is return on equity and b is retention ratio,

Retention ratio = 1 - payout ratio

= 1 - 0.62

= 0.38

Therefore,

Sustainable growth rate = [0.163(0.38)] / [1 – 0.163(0.38)]

Sustainable growth rate = 0.06194 / 1 - 0.06194

Sustainable growth rate = 0.06194 / 0.93806

Sustainable growth rate = .0660, or 6.60%

If the solution helped, please give it a thumbs up. Thank you.